1) ABC (american bread company) is a competitor in the product market but a monopolist in the labor market. ABC's production function is characterized as Q=30L-0.5L2. It is also known that the market demand and supply of bread was Q=1000-5P and Q=-20+250P respectevily. The market supply of labor is W=2L answer the following

A. derive the unconditional demand for labor for ABC. show all computations

b. suppose ABC does not exploit its monopsony powers ( acts as a competitor in the labor market) compute the equilibrium W and L .

c. suppose ABC does exploit its monopsony status what is the equilibrium W and L now. in addition compute economic rent and employer surplus

d. if the government set a minimum wage (full coverage) at $35 what is the employment ES(employer surplus) ER (economic rent) now. compare to when abc exploits its monopsony power who gains, who loses? what is the net gaim
Answer #1

Date I Page No. PeaRect Compt un ur duct Mashet then at eam thus at E9m8 ω PC50-L)= 30P L(P+3) Pertect Compt un_ labor Mkt t

o Co8t as total factroo (Laboo) Cost: wL12 MAC OTPC4 TR- PO P- 800-8l5 o08- 825 (800-8s15) ( 30-1) = 4L 80 eg m at

L* = 25.6

W= 2*L = 51.2